in 1992 life expectancy was 63.4. in 1998 it was 67.0
E= life expectanc in year t and let t = # of years since 1992
what is the linear function for this?
what would the life expectancy be if E(16)

Answer by richwmiller(17219)

your points are (0, 63.4) and (6,67)
use them to write an equation after finding m and b
y-y/x-x=m
then find b using y=mx+b
now that you have m and b plug them in y=mx+b where y=E(x)